# Cognitive Hierarchy in Day-to-day Network Flow Dynamics

### v1

@@ 7 Conclusions and Future Research
This paper developed a new modeling framework of day-to-day network flow dynamics that incorporates hierarchical cognitive levels (Camerer et al., 2004) to better capture travelers’ dynamic re-routing behaviors. The classical NTP and Logit dynamics, on behalf of the DUE and SUE cases, respectively, were extended into the framework. Considering heterogeneity in travelers’ strategic-reasoning levels, the proposed CH-NTP model significantly improved the goodness-of-fit for a recent virtual experiment, which traditional day-to-day models could not achieve. For both dynamics, stationary states and local stabilities near the user equilibria were theoretically analyzed and verified by numerical experiments.
@@
We find that in addition to converging to the classical user equilibria (DUE/SUE), the proposed dynamics may also converge to other equilibria, depending on the initial conditions and parameters. For example, with a large $\gamma$, different initial conditions lead to different equilibria, while under a small $\gamma$, various initial points tend to evolve to the same DUE tardily (see again Figures 5(a) and 5(b)). Moreover, when higher-step travelers can accurately predict lower-step ones’ switching propensities, thinking multiple steps ahead does not affect the local stability of the CH-NTP dynamic around the DUE, and the stability condition does not depend on the distribution of strategic-reasoning levels. In comparison, the stability condition for the CH-Logit dynamic largely depends on the proportion of heterogeneous travelers. Finally, we find that for both dynamics with $|K|=2$: (i) when 1-step travelers over-predict 0-step ones’ switching propensity, the stable region is shrunk, and when the over-prediction is severe (e.g., $\hat{\gamma}>\bar{\gamma}$ for the CH-NTP dynamic), the UE is always unstable; (ii) when 1-step travelers moderately under-predict 0-step ones’ re-routing tendency, the stable region is increased; and (iii) when the under-predictions are severe, the stable region is separated into two parts.
@@
Despite the fresh findings generated, our models still have limitations, which may direct future research. First, we select two widely-used day-to-day models from the literature as examples of the general modeling framework. Many more classical day-to-day models can be analyzed in the future (e.g., Smith, 1979b, 1983; Nagurney and Zhang, 1997; Jin, 2007; Xiao et al., 2019), including those that consider travelers’ learning behavior (Horowitz, 1984; Cantarella and Cascetta, 1995; Bie and Lo, 2010). We surmise that findings on local stability still generally hold when hierarchical cognitive behaviors are incorporated into these models, provided that the dynamic itself exhibits similar properties (e.g., negative definiteness of the Jacobian). Second, the virtual experiment data suggest that travelers’ re-routing tendencies become weaker through repeated game play. Further research can use time-varying parameters to capture such a phenomenon better. Third, future research can investigate how different measures, such as congestion tolling (Tan et al., 2015; Guo et al., 2016; Liu et al., 2017; Han et al., 2017), traffic control (Smith, 1979a; Köhler and Strehler, 2019), and information provision schemes (Liu and Szeto, 2020), can improve the system performance. Due to the multi-equilibria phenomenon, it would be interesting to drive the traffic system to the desired state via these measures. Han et al. (2017) showed the possibility of using congestion tolling to direct traffic evolution to the target stationary state from any initial states when multiple equilibria exist. Our current research efforts are aimed in this direction.
@@ Acknowledgement
The research was supported by the National Natural Science Foundation of China (Project No.’s 72201214 and 72025104) and the Sichuan Science and Technology Program (Project No. 2023NSFSC1035).
@@ Appendix A Definitions and theorems on local stability and block matrices
(Khalil, 2002; Antsaklis and Michel, 2006) Let $\boldsymbol{x}^{\dagger}$ be an equilibrium point of the system (A.1) and $\phi$ be continuously differentiable in a neighborhood of $\boldsymbol{x}^{\dagger}$ with radius $\epsilon$, $\mathcal{B}(\boldsymbol{x}^{\dagger},\epsilon)$. Let $J$ be the Jacobian evaluated at $\boldsymbol{x}^{\dagger}$, $J=\frac{\partial\phi(\boldsymbol{x})}{\partial\boldsymbol{x}}|_{\boldsymbol{x}=\boldsymbol{x}^{\dagger}}$. Then $\boldsymbol{x}^{\dagger}$ is:
@@
(adopted from Equation 0.8.5.13 in Horn and Johnson, 2012) For block matrices $S_{11}$, $S_{12}$, $S_{21}$ and $S_{22}$, if $S_{11}$ commutes with either $S_{12}$ or $S_{21}$, i.e., $S_{11}S_{12}=S_{12}S_{11}$ or $S_{11}S_{21}=S_{21}S_{11}$, then $\begin{vmatrix}S_{11}&S_{12}\\ S_{21}&S_{22}\end{vmatrix}=|S_{11}S_{22}-S_{21}S_{12}|$.
@@
Proof. (We give a proof for completeness because most proofs we have found deal with the case of $2\times 2$ matrices only.) If $S_{11}$ is replaced by a $m$-size identity matrix, we know by expanding the first column that
$\begin{vmatrix}I_{m}&S_{12}&S_{13}\\ \boldsymbol{0}&S_{22}&S_{23}\\ \boldsymbol{0}&\boldsymbol{0}&S_{33}\end{vmatrix}=\begin{vmatrix}I_{m-1}&S_{12}^{1-}&S_{13}^{1-}\\ \boldsymbol{0}&S_{22}&S_{23}\\ \boldsymbol{0}&\boldsymbol{0}&S_{33}\end{vmatrix}$ (A.5)
where $S_{12}^{1-}$ and $S_{13}^{1-}$ represent $S_{12}$ and $S_{13}$ with their first rows removed. Since when $m=1$ (A.5) reduces to $\begin{vmatrix}S_{22}&S_{23}\\ \boldsymbol{0}&S_{33}\end{vmatrix}$, by induction on $m$, we know the determinant on the LHS of (A.5) is also $\begin{vmatrix}S_{22}&S_{23}\\ \boldsymbol{0}&S_{33}\end{vmatrix}$, which equals $|S_{22}|\cdot|S_{33}|$, as per Theorem A.2. Denote $G$ as matrix $\begin{vmatrix}S_{22}&S_{23}\\ \boldsymbol{0}&S_{33}\end{vmatrix}$.
@@
Using the similar mathematical induction technique on RHS of (A.6), we know its determinant is simply $|S_{11}|$. Hence, (A.6) indicates that $|G^{-1}|\cdot|S|=|S_{11}|$. Since $|G^{-1}|=\frac{1}{|G|}$, $|S|=|S_{11}|\cdot\begin{vmatrix}S_{22}&S_{23}\\ \boldsymbol{0}&S_{33}\end{vmatrix}=|S_{11}|\cdot|S_{22}|\cdot|S_{33}|$.
The case of $G$ being not invertible is trivial since $G$ would have at least one non-pivot row, and so does $S$. Hence, $|S|=|S_{11}||S_{22}||S_{33}|=0$. ∎
@@ Appendix B Proof of Lemma 2
The projection for a single OD pair is very similar to the case of projection into a probability simplex (Chen and Ye, 2011). We first write the Lagrangian of the optimization problem:
$L(\boldsymbol{z},\boldsymbol{\mu},\boldsymbol{\tau})=\frac{1}{2}\left\lVert{\boldsymbol{x}-\boldsymbol{z}}\right\rVert^{2}-\boldsymbol{\mu}^{T}\boldsymbol{x}+\boldsymbol{\tau}^{T}(\Gamma\boldsymbol{x}-\eta\boldsymbol{d})$ (B.1)
where $\boldsymbol{\mu}\equiv(\mu_{rw},r\in R_{w},w\in W)^{T}$ and $\boldsymbol{\tau}\equiv(\tau_{w},w\in W)^{T}$. The optimal $\boldsymbol{x}^{\ast},\boldsymbol{\mu}^{\ast},\boldsymbol{\tau}^{\ast}$ must satisfy the following Karush-Kuhn-Tucker conditions:
@@
where $\odot$ is the Hadamard product operator. Let $E(\boldsymbol{z}_{w})=\{r\in R_{w}|x_{rw}^{\ast}>0\}$. From (B.4) we obtain that $\mu_{rw}^{\ast}=0,\forall r\in E(\boldsymbol{z}_{w}),\forall w$. Then from (B.2) we further know $x_{rw}^{\ast}=z_{rw}-\tau_{w}^{\ast},\forall r\in E(\boldsymbol{z}_{w}),\forall w$. Substituting it into the equality in (B.3) yields $\sum_{r\in E(\boldsymbol{z}_{w})}{(z_{rw}-\tau_{w}^{\ast})}=\eta d_{w},\forall w$. Hence $\tau_{w}^{\ast}=\frac{(\sum_{r\in E(\boldsymbol{z}_{w})}{z_{rw}})-\eta d_{w}}{|E(\boldsymbol{z}_{w})|}$. Using a similar deduction process, we can conclude that $\forall r\in R_{w}\setminus E(\boldsymbol{z}_{w})$, $z_{rw}\leq\tau_{w}^{\ast}$.
The above reasoning shows that the projection, for each $w$, has the form of:
@@
In a piecewise form, first note that this operator is undifferentiable only when the set $E(\boldsymbol{z}_{w})$ suddenly changes. The Jacobian of (B.5) can be expressed as:
@@
where $\mathds{1}_{w}$ is an indicator vector of size $|R_{w}|$ whose $r$-th entry is 1 if $r\in E(\boldsymbol{z}_{w})$ and 0 otherwise. When $|E(\boldsymbol{z}_{w})|$ = $|R_{w}|$, $Q_{w}[\boldsymbol{z}_{w}]$ degrades to $\bar{Q}_{w}=\text{Diag}(\boldsymbol{1}_{w})-\frac{\boldsymbol{1}_{w}\boldsymbol{1}_{w}^{T}}{|R_{w}|}$ where $\boldsymbol{1}_{w}$ is all-one vector of size $|R_{w}|$.
To show $Q_{w}[\boldsymbol{z}_{w}]$ is PSD for any $\boldsymbol{z}_{w}$, we only need to look at the non-zero entries in $\mathds{1}_{w}$, which form an $|E(\boldsymbol{z}_{w})|$-by-$|E(\boldsymbol{z}_{w})|$ matrix, $\tilde{Q}_{w}[\boldsymbol{z}_{w}]\equiv\text{Diag}(\boldsymbol{1}_{E_{w}})-\frac{\boldsymbol{1}_{E_{w}}\boldsymbol{1}_{E_{w}}^{T}}{|E(\boldsymbol{z}_{w})|}$, where $\boldsymbol{1}_{E_{w}}$ is all-one vector of size $|E(\boldsymbol{z}_{w})|$. Note that matrix $|E(\boldsymbol{z}_{w})|\cdot\tilde{Q}_{w}[\boldsymbol{z}_{w}]$ can be viewed as the Laplacian of a fully connected graph with the number of nodes being $|E(\boldsymbol{z}_{w})|+1$. It is well-known that the Laplacian of any graph is PSD. Hence, $|E(\boldsymbol{z}_{w})|\cdot\tilde{Q}_{w}[\boldsymbol{z}_{w}]$ is also PSD.
To show $(Q_{w}[\boldsymbol{z}_{w}])^{2}=Q_{w}[\boldsymbol{z}_{w}]$, we again only need to focus on the non-zero submatrix $\tilde{Q}_{w}[\boldsymbol{z}_{w}]$. It suffices to show $\tilde{Q}_{w}[\boldsymbol{z}_{w}](\tilde{Q}_{w}[\boldsymbol{z}_{w}]-\text{Diag}(\boldsymbol{1}_{E_{w}}))=0$. This is true because all the entries of $\tilde{Q}_{w}[\boldsymbol{z}_{w}]-\text{Diag}(\boldsymbol{1}_{E_{w}}$) are the same and entries of each column of $\tilde{Q}_{w}[\boldsymbol{z}_{w}]$ sum up to 0.
The above indicates that $Q_{w}[\boldsymbol{z}_{w}]$ is in fact an orthogonal projection matrix. So $\text{rank}(Q_{w}[\boldsymbol{z}_{w}])=tr(Q_{w}[\boldsymbol{z}_{w}])=|E(\boldsymbol{z}_{w})|-1$. Since rank is the number of nonzero eigenvalues, the algebraic multiplicity for $\lambda=1$ is also $|E(\boldsymbol{z}_{w})|-1$. Thanks to $(Q_{w}[\boldsymbol{z}_{w}])^{2}=Q_{w}[\boldsymbol{z}_{w}]$, every vector in the range of $Q_{w}[\boldsymbol{z}_{w}]$ is an eigenvector with eigenvalue 1 and thus the associated eigenvalue’s geometric multiplicity is $|E(\boldsymbol{z}_{w})|-1$. By the rank-nullity theorem, rank $(Q_{w}[\boldsymbol{z}_{w}]-I)$ is $|R_{w}|-|E(\boldsymbol{z}_{w})|+1$.

### v2

@@ 7 Conclusions and Future Research
This paper developed a new modeling framework of day-to-day network flow dynamics that incorporates cognitive hierarchy theory to better capture travelers’ dynamic re-routing behaviors. The classical NTP and Logit dynamics, on behalf of the DUE and SUE cases, respectively, were extended into the framework. By capturing heterogeneity in travelers’ strategic-reasoning levels, the proposed CH-NTP model significantly improved the goodness-of-fit for a recent virtual experiment, which traditional day-to-day models could not achieve.
@@
We find that in addition to converging to the user equilibria, the proposed dynamics may also converge to other equilibria, depending on the initial conditions and parameters. For example, our analysis of the CH-NTP dynamic reveals that non-DUE MPE emerge when travelers extensively anticipate others’ behaviors. Interestingly, non-DUE MPE can either improve or worsen system efficiency compared to DUE, depending on the proportion of strategic travelers and the cost functions. Moreover, when higher-step travelers can accurately predict lower-step ones’ switching propensities, thinking multiple steps ahead does not affect the local stability of the CH-NTP dynamic around the DUE. In comparison, the stability condition for the CH-Logit dynamic largely depends on the proportion of heterogeneous travelers. Finally, we find that for both dynamics with $|K|=2$: (i) when 1-step travelers over-predict 0-step ones’ switching propensity, the stable region is shrunk, and when the over-prediction is severe (e.g., $\hat{\gamma}>\bar{\gamma}$ for the CH-NTP dynamic), the UE is always unstable; (ii) when 1-step travelers moderately under-predict 0-step ones’ re-routing tendency, the stable region is increased; and (iii) when the under-predictions are severe, the stable region is separated into two parts.
@@
Despite the fresh findings generated, our models still have limitations, which may direct future research. First, we select two widely-used day-to-day models from the literature as examples of the general modeling framework. Many more day-to-day models can be analyzed in the future (e.g., Smith 1979b, 1983, Nagurney and Zhang 1997, Jin 2007, Xiao et al. 2019, Li et al. 2024). We surmise that findings on local stability still generally hold when hierarchical cognitive behaviors are incorporated into these models, provided that the dynamic itself exhibits regular properties (e.g., definiteness of the Jacobian). Second, it would also be interesting to apply the idea of cognitive hierarchy to investigate the departure time choice problem in the day-to-day setting (Guo, Yang, and Huang 2018, Jin 2020, 2021, Lamotte and Geroliminis 2021, Guo, Yang, and Huang 2023). Exploring how different levels of strategic thinking influence travelers’ departure time adjustment may reveal new equilibrium properties in day-to-day bottleneck dynamics. Third, the virtual experiment data suggest that travelers’ strategic thinking levels may evolve with experience, as participants appeared to adopt more uniformly random strategies after extended gameplay. Future research could incorporate time-varying parameters, particularly a time-decreasing prediction parameter, to better capture this behavioral adaptation. Fourth, future research can investigate how different measures, such as congestion tolling (Tan, Yang, and Guo 2015, Guo et al. 2016, Han et al. 2017, Wu, Yin, and Lynch 2024), traffic control (Smith 1979a, Köhler and Strehler 2019), and information provision schemes (Liu and Szeto 2020), can improve system performance. Due to the multi-equilibria phenomenon, it is be interesting to drive the traffic system to desired states via these measures.
@@ Acknowledgement
Correspondence concerning this article should be addressed to Feng Xiao. The research was supported by the National Natural Science Foundation of China (Project No.’s 72201214 and 72025104), the Sichuan Science and Technology Program (Project No. 2023NSFSC1035) and the Central Guidance for Local Science and Technology Development Fund Projects (Project No. 2024ZYD0108).
@@ Appendix A Definitions and theorems on local stability and block matrices
(Khalil 2002, Antsaklis and Michel 2006) Let $\boldsymbol{x}^{\dagger}$ be an equilibrium point of the system in Eq. (A.1) and $\phi$ be continuously differentiable in a neighborhood of $\boldsymbol{x}^{\dagger}$ with radius $\epsilon$, $\mathcal{B}(\boldsymbol{x}^{\dagger},\epsilon)$. Let $J$ be the Jacobian evaluated at $\boldsymbol{x}^{\dagger}$, $J=\frac{\partial\phi(\boldsymbol{x})}{\partial\boldsymbol{x}}|_{\boldsymbol{x}=\boldsymbol{x}^{\dagger}}$. Then $\boldsymbol{x}^{\dagger}$ is:
@@
(adapted from Equation 0.8.5.13 in Horn and Johnson 2012) For block matrices $S_{11}$, $S_{12}$, $S_{21}$ and $S_{22}$, if $S_{11}$ commutes with either $S_{12}$ or $S_{21}$, i.e., $S_{11}S_{12}=S_{12}S_{11}$ or $S_{11}S_{21}=S_{21}S_{11}$, then $\begin{vmatrix}S_{11}&S_{12}\\ S_{21}&S_{22}\end{vmatrix}=|S_{11}S_{22}-S_{21}S_{12}|$.
@@
Proof. (We give a proof for completeness because most proofs we have found deal with the case of $2\times 2$ matrices only.) If $S_{11}$ is replaced by a $m$-size identity matrix, expanding the first column yields:
$\begin{vmatrix}I_{m}&S_{12}&S_{13}\\ \boldsymbol{0}&S_{22}&S_{23}\\ \boldsymbol{0}&\boldsymbol{0}&S_{33}\end{vmatrix}=\begin{vmatrix}I_{m-1}&S_{12}^{1-}&S_{13}^{1-}\\ \boldsymbol{0}&S_{22}&S_{23}\\ \boldsymbol{0}&\boldsymbol{0}&S_{33}\end{vmatrix},$ (A.5)
where $S_{12}^{1-}$ and $S_{13}^{1-}$ represent $S_{12}$ and $S_{13}$ with their first rows removed. When $m=1$, Eq. (A.5) reduces to $\begin{vmatrix}S_{22}&S_{23}\\ \boldsymbol{0}&S_{33}\end{vmatrix}$. By induction on $m$, the determinant on the left-hand side of Eq. (A.5) equals $\begin{vmatrix}S_{22}&S_{23}\\ \boldsymbol{0}&S_{33}\end{vmatrix}$. According to Theorem A.2, this equals $|S_{22}|\cdot|S_{33}|$. Denote $G$ as matrix $\begin{vmatrix}S_{22}&S_{23}\\ \boldsymbol{0}&S_{33}\end{vmatrix}$.
@@
Using the similar mathematical induction technique on RHS of Eq. (A.6), we know its determinant is simply $|S_{11}|$. Hence, Eq. (A.6) indicates that $|G^{-1}|\cdot|S|=|S_{11}|$. Since $|G^{-1}|=\frac{1}{|G|}$, $|S|=|S_{11}|\cdot\begin{vmatrix}S_{22}&S_{23}\\ \boldsymbol{0}&S_{33}\end{vmatrix}=|S_{11}|\cdot|S_{22}|\cdot|S_{33}|$.
The case of $G$ being not invertible is trivial since $G$ would have at least one non-pivot row, and so does $S$. In this case, $|S|=|S_{11}||S_{22}||S_{33}|=0$. ∎
@@ Appendix B Proof of Lemma 2
Projecting a vector $\boldsymbol{z}$ onto $\Omega_{\eta}\equiv\{\boldsymbol{x}|\Gamma\boldsymbol{x}=\eta\boldsymbol{d},\boldsymbol{x}\geq 0\}$ requires solving the optimization problem $\operatorname*{arg\,min}_{y\in\Omega_{\eta}}\left\lVert{\boldsymbol{y}-\boldsymbol{z}}\right\rVert^{2}$. The Lagrangian of this optimization problem can be expressed as:
$L(\boldsymbol{z},\boldsymbol{\mu},\boldsymbol{\tau})=\frac{1}{2}\left\lVert{\boldsymbol{x}-\boldsymbol{z}}\right\rVert^{2}-\boldsymbol{\mu}^{T}\boldsymbol{x}+\boldsymbol{\tau}^{T}(\Gamma\boldsymbol{x}-\eta\boldsymbol{d}),$ (B.1)
where $\boldsymbol{\mu}\equiv(\mu_{rw},r\in R_{w},w\in W)^{T}$ and $\boldsymbol{\tau}\equiv(\tau_{w},w\in W)^{T}$ represent the Lagrangian multipliers. These multipliers correspond to the non-negative flow constraints and flow conservation constraints, respectively.
The optimal $\boldsymbol{x}^{\ast},\boldsymbol{\mu}^{\ast},\boldsymbol{\tau}^{\ast}$ must satisfy the following Karush-Kuhn-Tucker conditions:
@@
where $\odot$ is the Hadamard product operator. Let $E(\boldsymbol{z}_{w})=\{r\in R_{w}|x_{rw}^{\ast}>0\}$. From Eq. (B.4) we obtain that $\mu_{rw}^{\ast}=0,\forall r\in E(\boldsymbol{z}_{w}),\forall w$. Hence, from Eq. (B.2) we further know $x_{rw}^{\ast}=z_{rw}-\tau_{w}^{\ast},\forall r\in E(\boldsymbol{z}_{w}),\forall w$. Substituting $x_{rw}^{\ast}=z_{rw}-\tau_{w}^{\ast}$ into the equality in Eq. (B.3) yields $\sum_{r\in E(\boldsymbol{z}_{w})}{(z_{rw}-\tau_{w}^{\ast})}=\eta d_{w},\forall w$. Hence $\tau_{w}^{\ast}=\frac{(\sum_{r\in E(\boldsymbol{z}_{w})}{z_{rw}})-\eta d_{w}}{|E(\boldsymbol{z}_{w})|}$.
Let $\bar{E}(\boldsymbol{z}_{w})=\{r\in R_{w}|x_{rw}^{\ast}=0\}$. From Eq. (B.2) we know that $z_{rw}+\mu_{rw}^{\ast}-\tau_{w}^{\ast}=0,\forall r\in\bar{E}(\boldsymbol{z}_{w})$. Since $\mu_{rw}^{\ast}\geq 0,\forall r$, we can conclude that $z_{rw}\leq\tau_{w}^{\ast},\forall r\in\bar{E}(\boldsymbol{z}_{w})$.
The above two cases show that the projection, for each $w$, has the form of:
@@
In a piecewise form, first note that this operator is undifferentiable only when the set $E(\boldsymbol{z}_{w})$ suddenly changes. The Jacobian of Eq. (B.5) can be expressed as:
@@
where $\mathds{1}_{w}$ is an indicator vector of size $|R_{w}|$ whose $r$-th entry is 1 if $r\in E(\boldsymbol{z}_{w})$ and 0 otherwise. When $|E(\boldsymbol{z}_{w})|$ = $|R_{w}|$ (i.e., $x_{rw}^{\ast}>0,\forall r$), $Q_{w}[\boldsymbol{z}_{w}]$ simplifies to $\bar{Q}_{w}=\text{Diag}(\boldsymbol{1}_{w})-\frac{\boldsymbol{1}_{w}\boldsymbol{1}_{w}^{T}}{|R_{w}|}$ where $\boldsymbol{1}_{w}$ denotes the all-one vector of size $|R_{w}|$.
To prove that $Q_{w}[\boldsymbol{z}_{w}]$ is PSD for any $\boldsymbol{z}_{w}$, we first observe that the rows or columns corresponding to $\bar{E}(\boldsymbol{z}_{w})$ (where $\mathds{1}_{w}=0$) contribute only zeros to $Q_{w}[\boldsymbol{z}_{w}]$. Thus, for any vector $\boldsymbol{a}$, the quadratic form $\boldsymbol{a}^{T}Q_{w}[\boldsymbol{z}_{w}]\boldsymbol{a}$ simplifies to $\boldsymbol{a}_{E(\boldsymbol{z}_{w})}^{T}\tilde{Q}_{w}[\boldsymbol{z}_{w}]\boldsymbol{a}_{E(\boldsymbol{z}_{w})}$, where $\tilde{Q}_{w}[\boldsymbol{z}_{w}]$ is the submatrix of $Q_{w}[\boldsymbol{z}_{w}]$ formed by retaining only the rows and columns indexed by $E(\boldsymbol{z}_{w})$, and $\boldsymbol{a}_{E(\boldsymbol{z}_{w})}$ denotes the subvector of $\boldsymbol{a}$ indexed by $E(\boldsymbol{z}_{w})$. Therefore, it suffices to analyze $\tilde{Q}_{w}[\boldsymbol{z}_{w}]$, which forms an $|E(\boldsymbol{z}_{w})|$-by-$|E(\boldsymbol{z}_{w})|$ matrix, $\tilde{Q}_{w}[\boldsymbol{z}_{w}]\equiv\text{Diag}(\boldsymbol{1}_{E_{w}})-\frac{\boldsymbol{1}_{E_{w}}\boldsymbol{1}_{E_{w}}^{T}}{|E(\boldsymbol{z}_{w})|}$, where $\boldsymbol{1}_{E_{w}}$ is the all-one vector of size $|E(\boldsymbol{z}_{w})|$. Note that matrix $|E(\boldsymbol{z}_{w})|\cdot\tilde{Q}_{w}[\boldsymbol{z}_{w}]$ can be viewed as the Laplacian of a fully connected graph with the number of nodes being $|E(\boldsymbol{z}_{w})|+1$. It is well-known that the Laplacian of any graph is PSD. Therefore, $\tilde{Q}_{w}[\boldsymbol{z}_{w}]$ is also PSD, since scaling by $\frac{1}{|E(\boldsymbol{z}_{w})|}$ preserves the PSD.
To show $(Q_{w}[\boldsymbol{z}_{w}])^{2}=Q_{w}[\boldsymbol{z}_{w}]$ (i.e., idempotent), we again only need to focus on the non-zero submatrix $\tilde{Q}_{w}[\boldsymbol{z}_{w}]$. It suffices to show $\tilde{Q}_{w}[\boldsymbol{z}_{w}](\tilde{Q}_{w}[\boldsymbol{z}_{w}]-\text{Diag}(\boldsymbol{1}_{E_{w}}))=0$. This holds true because all the entries of $\tilde{Q}_{w}[\boldsymbol{z}_{w}]-\text{Diag}(\boldsymbol{1}_{E_{w}}$) are the same and entries of each column of $\tilde{Q}_{w}[\boldsymbol{z}_{w}]$ sum up to 0. To show $Q_{w}[\boldsymbol{z}_{w}]$ is symmetric, we note that $\text{Diag}(\mathds{1}_{w})$ and $\mathds{1}_{w}\mathds{1}_{w}^{T}$ are symmetric, so their difference $Q_{w}[\boldsymbol{z}_{w}]$ is symmetric.
The idempotence and symmetry of $Q_{w}[\boldsymbol{z}_{w}]$ tell that the matrix is an orthogonal projection matrix. So $\text{rank}(Q_{w}[\boldsymbol{z}_{w}])=tr(Q_{w}[\boldsymbol{z}_{w}])=|E(\boldsymbol{z}_{w})|-1$. Since rank is the number of nonzero eigenvalues, the algebraic multiplicity for $\lambda=1$ is also $|E(\boldsymbol{z}_{w})|-1$. Thanks to $(Q_{w}[\boldsymbol{z}_{w}])^{2}=Q_{w}[\boldsymbol{z}_{w}]$, every vector in the range of $Q_{w}[\boldsymbol{z}_{w}]$ is an eigenvector with eigenvalue 1 and thus the associated eigenvalue’s geometric multiplicity is $|E(\boldsymbol{z}_{w})|-1$. By the rank-nullity theorem, rank $(Q_{w}[\boldsymbol{z}_{w}]-I)$ is $|R_{w}|-|E(\boldsymbol{z}_{w})|+1$.